# SpotFormer: Multi-Scale Spatio-Temporal Transformer for Facial Expression Spotting

### v2

@@ 2.4 Contrastive learning
In recent years, contrastive learning has proven to be effective in various domains, including computer vision and natural language processing. The objective of contrastive learning in unsupervised learning [chen2020simple] is to acquire discriminative representations by maximizing the similarity between similar instances (positive pairs generated using data augmentation) while minimizing the similarity between dissimilar instances (negative pairs). This approach produces meaningful visual representations applicable to tasks such as image recognition, representation learning, and semantic understanding. Based on the idea that positive pairs can also be selected by ground-truth labels, Khosla et al. [khosla2020supervised] introduced the supervised contrastive loss, showcasing its potential to enhance supervised tasks by incorporating labeled information during the training process. Supervised contrastive learning efficiently enlarges the domain discrepancy, leading to an enhanced extraction of discriminative feature representations. Our method focuses on recognizing the boundary between different types of expressions at the frame level. To achieve this objective, we establish contrasts between MaE and ME frames, as well as between ME frames and neutral frames. This approach enables our model to acquire more discriminative feature representations, ultimately reducing the misclassification rate.
@@ 3 Proposed Expression Spotting Framework
Figure 4: Overview of the proposed SpotFormer. (a) The network structure of SpotFormer. The dotted gray box provides an intuitive illustration of the spatio-temporal attention block. A spatial graph attention layer models the spatial relationships within the graph of each frame, while a temporal node attention layer captures temporal variations of each graph node across frames. Subsequently, spatial and temporal downsampling are performed using FLGP and a learnable temporal downsampling layer, respectively. (b) The scale change between different facial graph structures through FLGP. Each facial muscle is represented by a distinct color (e.g., orange represents the lips). In the first spatial downsampling, ROIs belonging to the same facial muscle are aggregated into single nodes via FLGP to construct graph #2. In the second spatial downsampling, all aggregated nodes are further merged into a single node via FLGP to form graph #3.
(c) The illustration of the spatial graph attention layer and temporal node attention layer.
@@ 3.1 Data pre-processing
Assuming that a video $V=(v_{i})_{i=1}^{N}$ with $N$ frames, we initially calculate the SW-MRO features. First, by setting a window length $w$, we pad the beginning and end of the video with $\lfloor\frac{w}{2}\rfloor$ repetitions of the first and last frames, respectively. This ensures that the original first and last frames can be the center frames of the first and last sliding windows, respectively. Then, we divide the entire video into $N$ overlapped clips $C=(c_{i})_{i=1}^{N}$ with the window length $w$ and sliding stride 1, where $c_{i}$ involves $v_{i}$ as the center frame. In this paper, we aim to achieve accurate frame-level probability estimation based on compact sliding windows. Specifically, we utilize all the temporal information of $c_{i}$ and magnify the motion information existing in $c_{i}$ to estimate the probability map of $v_{i}$. Next, we will introduce the process of obtaining multi-temporal-resolution optical flow features in each compact sliding window.
@@
For each clip $c_{i}$, we initially detect the 68 facial keypoints using a pre-trained MobileFaceNet [PFL] in the first frame $c_{i}^{1}$. These keypoints will subsequently be used to generate the facial bounding box, extract the nose tip area for facial alignment, and extract ROIs. Specifically, we detect the facial bounding box in the first frame $c_{i}^{1}$. In each subsequent frame $c_{i}^{s}$, where $s=2,3,\ldots,w$, we initialize the facial bounding box of $c_{i}^{s}$ with the one from $c_{i}^{1}$. Given that the nose tip area remains stationary during facial expressions [yuhong2021research], we compute the optical flow of the nose tip area, as illustrated in Fig. 3, to represent the global head movement from $c_{i}^{1}$ to $c_{i}^{s}$. To achieve this, we employ the Farneback algorithm [farneback2003two] to compute MDMO optical flow [liu2015main] of the nose tip area $o_{i}^{s,\mathrm{nose}}\in\mathbb{R}^{2}$ from $c_{i}^{1}$ to $c_{i}^{s}$:
$o^{s,\mathrm{nose}}_{i}=\frac{1}{m^{\mathrm{nose}}_{i}\times n^{\mathrm{nose}}_{i}}\sum_{(x,y)\in M^{\mathrm{nose}}_{i}}\mathrm{OF}(c_{i}^{1},c_{i}^{s})[x,y],$ (1)
@@
where $m^{\mathrm{nose}}_{i}$ and $n^{\mathrm{nose}}_{i}$ represent the height and width of the extracted nose tip area $M^{\mathrm{nose}}_{i}$, and $\mathrm{OF}(\cdot,\cdot)[x,y]$ denotes the value of MDMO optical flow features at $(x,y)$ coordinates. Consequently, $o^{s,\mathrm{nose}}_{i}$ contains the horizontal and vertical movements of the head from $c_{i}^{1}$ to $c_{i}^{s}$. Then, we translate the facial bounding box of frame $c_{i}^{s}$ by adding $o^{s,\mathrm{nose}}_{i}$ to complete facial alignment.
@@
Based on the discussion in Sections 1 and 2, we selectively extract $R$ ROIs [zhao2022rethinking], where facial expressions happen most frequently, rather than processing entire images. This choice can effectively mitigate the impact of noise and irrelevant facial movements. Then, we compute the optical flow for the selected $R$ ROIs between the first frame $c_{i}^{1}$ and frame $c_{i}^{s}$ to obtain MDMO optical flow features $o_{i}^{s}=[o_{i}^{s,r}]_{r=1}^{R}\in\mathbb{R}^{R\times 2}$. The computation is similar to Eq. (1). Afterwards, we concatenate the optical flow $o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{w}]$ to construct multi-temporal-resolution optical flow features $o_{i}\in\mathbb{R}^{w\times R\times 2}$ for the clip $c_{i}$, where $o_{i}^{1}=\bm{0}$. As a result of the data pre-processing, we concatenate the multi-temporal-resolution optical flow from all clips and obtain the sliding window-based multi-temporal-resolution optical flow (SW-MRO) features $O=(o_{i})_{i=1}^{N}\in\mathbb{R}^{N\times w\times R\times 2}$ for the entire input video.
@@ 3.2.2 Temporal positional embedding
For the temporal dimension, following [dosovitskiy2021an], we set tokens with learnable parameters, denoted as $\delta_{\mathrm{t}}\in\mathbb{R}^{w\times D}$, to retain temporal positional information. Note that all graph patches in a single frame share the same temporal positional embedding.
@@ 3.2.3 Spatial facial embedding
Unlike other graph-based tasks where there are predefined edge connections among graph nodes, facial expression analysis involves many possible edge connections among different facial muscles. To mitigate the impact of complex and diverse edge connections, SpotFormer constructs parameterized associations between nodes based on facial muscle characteristics, unlike traditional GCNs that rely on pre-defined edge connections. Inspired by [zhang2021STST], we introduce learnable facial part tokens to the graph nodes. Specifically, we set three tokens with learnable parameters $\delta_{\mathrm{s}}\in\mathbb{R}^{3\times D}$ for three asymmetric facial parts (left eyebrow, right eyebrow, and mouth) since MEs are asymmetric. Graph nodes belonging to the same facial part share the same facial part token. This method enables the model to understand the overall spatial facial structure without the requirement of exploring complex pre-defined edge connections.
@@ 3.2.5 Multi-scale spatio-temporal feature learning
Multi-scale learning has shown powerful performance in image processing [lin2017feature, liu2021swin] and video understanding [Fan_2021_ICCV, Shao_2023_ICCV]. It is also significant in facial expression analysis since it enables the model to extract both coarse and fine feature representations across different scales, enhancing its ability to capture comprehensive facial structures and temporal dynamics.
SpotFormer extracts multi-scale spatio-temporal features, enabling more comprehensive learning of motion patterns from low-level to high-level. Specifically, to extract multi-scale temporal features, we use 1D CNNs with a stride of 2 for temporal downsampling, as proposed in ActionFormer [zhang2022actionformer].
However, for spatial downsampling, applying pooling operations, which are generally used for downsampling images, to graph-structured data presents challenges because it is a type of non-Euclidean structured data. To address this issue, inspired by Xu et al. [xu2021graph], we introduce FLGP, specifically designed for extracting multi-scale facial graph features. In practice, we design three scales of facial structures. The designed scales and the scale change achieved through FLGP are illustrated in Fig. 4 (b). During each FLGP operation, the facial graph is downsampled by aggregating features from multiple nodes using the max pooling operation.
@@ 3.2.7 Frame-level probability estimation
Finally, after reducing the spatial graph to a single node and temporal dimension to 1 through SpotFormer, the remaining single node comprehensively aggregates spatio-temporal information within a compact sliding window. We then use a fully-connected (FC) layer to generate the probability map $p_{i}=\{p_{i}^{\mathrm{onset}}$, $p_{i}^{\mathrm{apex}}$, $p_{i}^{\mathrm{offset}}$, $p_{i}^{\mathrm{exp}}$, $p_{i}^{\mathrm{neut}}$} corresponding to the frame $v_{i}$. This map contains the probabilities of $v_{i}$ being an onset frame, apex frame, offset frame, expression frame, or neutral frame. Additionally, each component in $p_{i}$ incorporates two probabilities for ME spotting and MaE spotting, respectively. Specifically, $p_{i}^{\mathrm{onset}}=\{p_{i}^{\mathrm{me},\mathrm{onset}},p_{i}^{\mathrm{mae},\mathrm{onset}}\}$, $p_{i}^{\mathrm{apex}}=\{p_{i}^{\mathrm{me},\mathrm{apex}},p_{i}^{\mathrm{mae},\mathrm{apex}}\}$, $p_{i}^{\mathrm{offset}}=\{p_{i}^{\mathrm{me},\mathrm{offset}},p_{i}^{\mathrm{mae},\mathrm{offset}}\}$, $p_{i}^{\mathrm{exp}}=\{p_{i}^{\mathrm{me},\mathrm{exp}},p_{i}^{\mathrm{mae},\mathrm{exp}}\}$, $p_{i}^{\mathrm{neut}}=\{p_{i}^{\mathrm{me},\mathrm{neut}},p_{i}^{\mathrm{mae},\mathrm{neut}}\}$. Therefore, the model outputs ten probabilities for each frame. The optimization tasks are divided into two groups for MaE spotting and ME spotting, respectively. Each group contains two binary classification tasks and a three-class classification task for different types of frames, following the optimization method outlined in [leng2022abpn]. Specifically, the first binary classification task determines whether the current frame is an apex frame, and the second determines whether it is an expression frame. The three-class classification task is used to classify each frame as an onset frame, an offset frame, or neither. Focal-loss [lin2017focal] is employed to optimize our model, which can be expressed as:
@@ 3.3 Supervised contrastive learning
To overcome this issue, we introduce supervised contrastive learning [khosla2020supervised] to enhance our model’s discriminative feature learning, aiming to improve its ability to recognize the classification boundaries between different types of expressions at the frame level. Specifically, we utilize the feature representation before the last classification FC layer as the feature representation for each frame. The supervised contrastive loss is then employed to minimize the distance between feature representations of the same expression class while simultaneously pushing apart feature representations of different expression classes. The expression type label $\widetilde{y}_{i}$ of the frame $v_{i}$ for the $i$-th sliding window $c_{i}$ is used as the supervision information for the supervised contrastive loss. This means that each frame is labeled as a MaE frame, ME frame, or neutral frame. Let $I$ denote a set of samples in a batch, and the loss function can be expressed as:
@@
Traditional methods MDMD[he2020spotting] 0.0629 0.0364 0.0445 0.1196 0.0082 0.0376
Optical Strain[gan2020optical] - - - 0.1436 0.0098 0.0448
Zhang et al.[zhang2020spatio] 0.0725 0.1331 0.0999 0.2131 0.0547 0.1403
He [yuhong2021research] 0.4149 0.2162 0.3638 0.3782 0.1965 0.3436
Zhao et al.[zhao2022rethinking] - - 0.3863 - - 0.4030
Wang et al.[wangunique] 0.3724 0.2866 0.3419 0.5061 0.2614 0.4558
Deep-learning methods Verburg[verburg2019micro] - 0.0821 - - - -
LBCNN[pan2020local] - - 0.0813 - - 0.0595
MESNet[wang2021mesnet] - 0.0880 - - 0.0360 -
SOFTNet[liong2021shallow] 0.2169 0.1520 0.1881 0.2410 0.1173 0.2022
3D-CNN[yap20223d] 0.1595 0.0466 0.1084 0.2145 0.0714 0.1675
Concat-CNN[yang2021facial] 0.3553 0.1155 0.2736 0.2505 0.0153 0.2019
LSSNet[yu2021lssnet] 0.2810 0.1310 0.2380 0.3770 0.0420 0.3250
MTSN[liong2022mtsn] 0.3459 0.0878 0.2867 0.4104 0.0808 0.3620
ABPN[leng2022abpn] 0.3349 0.1689 0.2908 0.3357 0.1590 0.3117
AUW-GCN[yin2023aware] 0.4293 0.1984 0.3728 0.4235 0.1538 0.3834
LGSNet[yu2023lgsnet] - - 0.3880 - - 0.4360
MULT[GUO2023146] - 0.2770 - - 0.1373 -

### v3

@@ 2.4 Contrastive learning
In recent years, contrastive learning has proven to be effective in various domains, including computer vision and natural language processing. The objective of contrastive learning in unsupervised learning [7] is to acquire discriminative representations by maximizing the similarity between similar instances (positive pairs generated using data augmentation) while minimizing the similarity between dissimilar instances (negative pairs). This approach produces meaningful visual representations applicable to tasks such as image recognition, representation learning, and semantic understanding. Based on the idea that positive pairs can also be selected by ground-truth labels, Khosla et al. [25] introduced the supervised contrastive loss, showcasing its potential to enhance supervised tasks by incorporating labeled information during the training process. Supervised contrastive learning efficiently enlarges the domain discrepancy, leading to an enhanced extraction of discriminative feature representations. Our method focuses on recognizing the boundary between different types of expressions at the frame level. To achieve this objective, we establish contrasts between MaE and ME frames, as well as between ME frames and neutral frames. This approach enables our model to acquire more discriminative feature representations, ultimately reducing the misclassification rate.
@@ 3 Proposed Expression Spotting Framework
Figure 4: Overview of the proposed SpotFormer. (a) The network structure of SpotFormer. The dotted gray box provides an intuitive illustration of the spatio-temporal attention block. A spatial graph attention layer models the spatial relationships within the graph of each frame, while a temporal node attention layer captures temporal variations of each graph node across frames. Subsequently, spatial and temporal downsampling are performed using FLGP and a learnable temporal downsampling layer, respectively. (b) The scale change between different facial graph structures through FLGP. Each facial muscle is represented by a distinct color (e.g., orange represents the lips). In the first spatial downsampling, ROIs belonging to the same facial muscle are aggregated into single nodes via FLGP to construct graph #2. In the second spatial downsampling, all aggregated nodes are further merged into a single node via FLGP to form graph #3. (c) The illustration of the spatial graph attention layer and temporal node attention layer.
@@ 3.1 Data pre-processing
Assuming that a video $V=(v_{i})_{i=1}^{N}$ with $N$ frames, we initially calculate the CSW-MRO features. First, by setting a window length $w$, we pad the beginning and end of the video with $\lfloor\frac{w}{2}\rfloor$ repetitions of the first and last frames, respectively. This ensures that the original first and last frames can be the center frames of the first and last sliding windows, respectively. Then, we divide the entire video into $N$ overlapped clips $C=(c_{i})_{i=1}^{N}$ with the window length $w$ and sliding stride 1, where $c_{i}$ involves $v_{i}$ as the center frame. In this paper, we aim to achieve accurate frame-level probability estimation based on compact sliding windows. Specifically, we utilize all the temporal information of $c_{i}$ and magnify the motion information existing in $c_{i}$ to estimate the probability map of $v_{i}$. Next, we will introduce the process of obtaining multi-temporal-resolution optical flow features in each compact sliding window.
@@
For each clip $c_{i}$, we initially detect the 68 facial keypoints using a pre-trained MobileFaceNet [6] in the first frame $c_{i}^{1}$. These keypoints will subsequently be used to generate the facial bounding box, extract the nose tip area for facial alignment, and extract ROIs. Specifically, we detect the facial bounding box in the first frame $c_{i}^{1}$. In each subsequent frame $c_{i}^{s}$, where $s=2,3,\ldots,w$, we initialize the facial bounding box of $c_{i}^{s}$ with the one from $c_{i}^{1}$. Given that the nose tip area remains stationary during facial expressions [77], we compute the optical flow of the nose tip area, as illustrated in Fig. 3, to represent the global head movement from $c_{i}^{1}$ to $c_{i}^{s}$. To achieve this, we employ the Farneback algorithm [17] to compute MDMO optical flow [39] of the nose tip region from $c_{i}^{1}$ to $c_{i}^{s}$. The resulting feature $o_{i}^{s,\mathrm{nose}}\in\mathbb{R}^{2}$ is obtained by averaging the optical flow vectors within the extracted nose region $M_{i}^{\mathrm{nose}}$ that align with the main direction of optical flow:
$o_{i}^{s,\mathrm{nose}}=\frac{1}{n_{i}^{\mathrm{nose}}}\sum_{(x,y)\in M_{i}^{\mathrm{nose}}}\mathrm{MDMO}(c_{i}^{1},c_{i}^{s})[x,y],$ (1)
where $n_{i}^{\mathrm{nose}}$ denotes the number of pixels in $M_{i}^{\mathrm{nose}}$ whose motion directions are consistent with the main optical flow direction, and $\mathrm{MDMO}(c_{i}^{1},c_{i}^{s})[x,y]$ represents the MDMO optical flow vector at location $(x,y)$. Optical flow vectors that do not follow the main direction are suppressed by setting them to $\vec{0}$.
@@
Consequently, $o^{s,\mathrm{nose}}_{i}$ contains the horizontal and vertical movements of the head from $c_{i}^{1}$ to $c_{i}^{s}$. Then, we translate the facial bounding box of frame $c_{i}^{s}$ by adding $o^{s,\mathrm{nose}}_{i}$ to complete facial alignment.
@@
Based on the discussion in Sections 1 and 2, we selectively extract $R$ ROIs [84], where facial expressions happen most frequently, rather than processing entire images. This choice can effectively mitigate the impact of noise and irrelevant facial movements. Then, we compute the optical flow for the selected $R$ ROIs between the first frame $c_{i}^{1}$ and frame $c_{i}^{s}$ to obtain MDMO optical flow features $o_{i}^{s}=[o_{i}^{s,r}]_{r=1}^{R}\in\mathbb{R}^{R\times 2}$. The computation is similar to Eq. (1). Afterwards, we concatenate the optical flow $o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{w}]$ to construct multi-temporal-resolution optical flow features $o_{i}\in\mathbb{R}^{w\times R\times 2}$ for the clip $c_{i}$, where $o_{i}^{1}=\bm{0}$. As a result of the data pre-processing, we concatenate the multi-temporal-resolution optical flow from all clips and obtain the compact sliding-window-based multi-temporal-resolution optical flow (CSW-MRO) features $O=(o_{i})_{i=1}^{N}\in\mathbb{R}^{N\times w\times R\times 2}$ for the entire input video.
It is important to emphasize that the proposed CSW-MRO feature is not a simple reuse of existing techniques, but a task-driven redesign tailored for FES. Existing methods adopt diverse strategies for optical flow computation. In many deep learning methods that compute optical flow for temporal modeling, it is often computed between adjacent frames, which provides robustness to head movement but may limit the magnification of motion information for detecting MEs. Moreover, the temporal range for optical flow modeling is either implicitly determined by the network architecture (e.g., the receptive fields of designed CNNs) or explicitly specified as a predefined hyper-parameter (e.g., in pretrained I3D-based methods during feature extraction), without explicit investigation into its optimality for distinguishing MaEs and MEs [47, 57, 75, 27, 71, 76, 21, 32, 67]. In contrast, some recent traditional methods explore computing optical flow between a reference frame and subsequent frames over long temporal windows, which can enhance motion magnitude but may be more sensitive to head movements [77, 73, 84, 48, 72, 66].
We identify these limitations and combine the complementary advantages of both strategies by computing multi-temporal-resolution optical flow within compact sliding windows, which explicitly balances motion magnification and robustness to head movement. In addition, we further investigate the impact of window length and provide empirical insights into its role in distinguishing MaEs and MEs (see Section 4.4.2), thereby enabling accurate frame-level probability estimation for FES.
@@ 3.2.2 Temporal positional embedding
For the temporal dimension, following [12], we set tokens with learnable parameters, denoted as $\delta_{\mathrm{t}}\in\mathbb{R}^{w\times D}$, to retain temporal positional information. Note that all graph patches in a single frame share the same temporal positional embedding.
@@ 3.2.3 Spatial facial embedding
Unlike other graph-based tasks where there are predefined edge connections among graph nodes, facial expression analysis involves many possible edge connections among different facial muscles. To mitigate the impact of complex and diverse edge connections, SpotFormer constructs parameterized associations between nodes based on facial muscle characteristics, unlike traditional GCNs that rely on pre-defined edge connections. Inspired by [81], we introduce learnable facial part tokens to the graph nodes. Specifically, we set three tokens with learnable parameters $\delta_{\mathrm{s}}\in\mathbb{R}^{3\times D}$ for three asymmetric facial parts (left eyebrow, right eyebrow, and mouth) since MEs are asymmetric. Graph nodes belonging to the same facial part share the same facial part token. This method enables the model to understand the overall spatial facial structure without the requirement of exploring complex pre-defined edge connections.
@@ 3.2.5 Multi-scale spatio-temporal feature learning
Multi-scale learning has shown powerful performance in image processing [35, 40] and video understanding [16, 52]. It is also significant in facial expression analysis since it enables the model to extract both coarse and fine feature representations across different scales, enhancing its ability to capture comprehensive facial structures and temporal dynamics.
SpotFormer extracts multi-scale spatio-temporal features, enabling more comprehensive learning of motion patterns from low-level to high-level. Specifically, to extract multi-scale temporal features, we use 1D CNNs with a stride of 2 for temporal downsampling, as proposed in ActionFormer [78].
However, for spatial downsampling, applying pooling operations, which are generally used for downsampling images, to graph-structured data presents challenges because it is a type of non-Euclidean structured data. To address this issue, inspired by Xu et al. [63], we introduce FLGP, specifically designed for extracting multi-scale facial graph features. In practice, we design three scales of facial structures. The designed scales and the scale change achieved through FLGP are illustrated in Fig. 4 (b). During each FLGP operation, the facial graph is downsampled by aggregating features from multiple nodes using the max pooling operation.
@@ 3.2.7 Frame-level probability estimation
Finally, after reducing the spatial graph to a single node and temporal dimension to 1 through SpotFormer, the remaining single node comprehensively aggregates spatio-temporal information within a compact sliding window. We then use a fully-connected (FC) layer to generate the probability map $p_{i}=\{p_{i}^{\mathrm{apex}}$, $p_{i}^{\mathrm{exp}}$, $p_{i}^{\mathrm{onset}}$, $p_{i}^{\mathrm{offset}}$, $p_{i}^{\mathrm{nb}}$} corresponding to the frame $v_{i}$. This map contains the probabilities of $v_{i}$ being an apex frame, expression frame, onset frame, offset frame, or non-boundary frame. Additionally, each component in $p_{i}$ incorporates two probabilities for ME spotting and MaE spotting, respectively. Specifically, $p_{i}^{\mathrm{apex}}=\{p_{i}^{\mathrm{me},\mathrm{apex}},p_{i}^{\mathrm{mae},\mathrm{apex}}\}$, $p_{i}^{\mathrm{exp}}=\{p_{i}^{\mathrm{me},\mathrm{exp}},p_{i}^{\mathrm{mae},\mathrm{exp}}\}$, $p_{i}^{\mathrm{onset}}=\{p_{i}^{\mathrm{me},\mathrm{onset}},p_{i}^{\mathrm{mae},\mathrm{onset}}\}$, $p_{i}^{\mathrm{offset}}=\{p_{i}^{\mathrm{me},\mathrm{offset}},p_{i}^{\mathrm{mae},\mathrm{offset}}\}$, $p_{i}^{\mathrm{nb}}=\{p_{i}^{\mathrm{me},\mathrm{nb}},p_{i}^{\mathrm{mae},\mathrm{nb}}\}$. Therefore, the model outputs ten probabilities for each frame. The optimization tasks are divided into two groups for MaE spotting and ME spotting, respectively. Each group contains two binary classification tasks and a three-class classification task for different types of frames, following the optimization method outlined in [27]. Specifically, the first binary classification task uses $p_{i}^{\mathrm{apex}}$ to determine whether the current frame is an apex frame. The second binary classification task uses $p_{i}^{\mathrm{exp}}$ to determine whether it is an expression or a neutral frame. The three-class classification task uses $p_{i}^{\mathrm{onset}}$, $p_{i}^{\mathrm{offset}}$, and $p_{i}^{\mathrm{nb}}$ to classify the frame as onset, offset, or non-boundary.
Focal-loss [36] is employed to optimize our model, which can be expressed as:
@@ 3.3 Supervised contrastive learning
To overcome this issue, we introduce supervised contrastive learning [25] to enhance our model’s discriminative feature learning, aiming to improve its ability to recognize the classification boundaries between different types of expressions at the frame level. Specifically, we utilize the feature representation before the last classification FC layer as the feature representation for each frame. The supervised contrastive loss is then employed to minimize the distance between feature representations of the same expression class while simultaneously pushing apart feature representations of different expression classes. The expression type label $\widetilde{y}_{i}$ of the frame $v_{i}$ for the $i$-th sliding window $c_{i}$ is used as the supervision information for the supervised contrastive loss. This means that each frame is labeled as a MaE frame, ME frame, or neutral frame. Let $I$ denote a set of samples in a batch, and the loss function can be expressed as:
@@
Traditional methods MDMD[23] 0.0629 0.0364 0.0445 0.1196 0.0082 0.0376
Optical Strain[19] - - - 0.1436 0.0098 0.0448
Zhang et al.[80] 0.0725 0.1331 0.0999 0.2131 0.0547 0.1403
He [77] 0.4149 0.2162 0.3638 0.3782 0.1965 0.3436
Zhao et al.[84] - - 0.3863 - - 0.4030
Wang et al.[56] 0.3724 0.2866 0.3419 0.5061 0.2614 0.4558
Deep-learning methods Verburg[55] - 0.0821 - - - -
LBCNN[46] - - 0.0813 - - 0.0595
MESNet[57] - 0.0880 - - 0.0360 -
SOFTNet[38] 0.2169 0.1520 0.1881 0.2410 0.1173 0.2022
3D-CNN[70] 0.1595 0.0466 0.1084 0.2145 0.0714 0.1675
Concat-CNN[65] 0.3553 0.1155 0.2736 0.2505 0.0153 0.2019
LSSNet[75] 0.2810 0.1310 0.2380 0.3770 0.0420 0.3250
MTSN[37] 0.3459 0.0878 0.2867 0.4104 0.0808 0.3620
ABPN[27] 0.3349 0.1689 0.2908 0.3357 0.1590 0.3117
AUW-GCN[71] 0.4293 0.1984 0.3728 0.4235 0.1538 0.3834
LGSNet[76] - - 0.3880 - - 0.4360
MULT[20] - 0.2770 - - 0.1373 -